# Artificial Human Intelligence: The role of Humans in the Development of Next Generation AI

### v1

@@ Abstract
Human intelligence, the most evident and accessible form of source of reasoning, hosted by biological hardware, has evolved and been refined over thousands of years, positioning itself today to create new artificial forms and preparing to self–design their evolutionary path forward. Beginning with the advent of foundation models, the rate at which human and artificial intelligence interact with each other has surpassed any anticipated quantitative figures. The close engagement led to both bits of intelligence to be impacted in various ways, which naturally resulted in complex confluences that warrant close scrutiny. In the sequel, we shall explore the interplay between human and machine intelligence, focusing on the crucial role humans play in developing ethical, responsible, and robust intelligent systems. We slightly delve into interesting aspects of implementation inspired by the mechanisms underlying neuroscience and human cognition. Additionally, we propose future perspectives, capitalizing on the advantages of symbiotic designs to suggest a human-centered direction for next-generation AI development. We finalize this evolving document with a few thoughts and open questions yet to be addressed by the broader community.
Keywords: Human-inspired AI, neuroscience, human vision, computer vision, cognition, neural networks, biological plausibility, CNN, KAN, Transformers.
1 Introduction
@@
The complexity of the intelligence problem, whose definition is yet to be established, is tempting to adapt human model in search of any type of intelligent behavior. Paradoxically, any conceivable form of artificial intelligence will inevitably originate either by mimicking human form or be a product of human intellect, at least in the near term, unless artificially generated forms begin building up newer forms of intelligences. Consequently, bias might be an intrinsic part of its initial design. However, implementation constraints, task variability, and interactions with other intelligent forms could guide humans in creating general intelligence to collaborate with [2]. Moreover, this collaboration could guide humanity to discover new inventions or new cognitive technologies, which might potentially expand the range of human thought. This virtuous feedback could eventually lead to artificially generated human-like intelligence, a byproduct of symbiotic evolution of digital and biological computational models.
@@
In this brief exposition, our objective is to investigate the strategic placement of human–centered inquiries in the advancement of contemporary AI systems despite the recent remarkable (not necessarily human–centric) developments in machine learning and computer vision fields [3]. Considering the strategies of machine and human intelligences in completing various tasks and their complementarity in behaviour, a potential fusion of both emerge as a plausible option that might soon become an applicable reality. Since human-centricity seems to be unavoidable in any type of AI-related research, we have decided to label this specific fusion of intelligences and all the interactions within as “artificial human intelligence” and explicate various intriguing topics of discourse for the mutually beneficial association of the two (seemingly) mechanistically distinct notions. Naturally, the exploration of both forms of intelligence necessitates addressing numerous intriguing questions. These include early human brain developments and their connections to neuroscience, as well as biological hardware dependencies, efficiency of implementations and the potential long-term effects on both human and machine intelligences as both intelligent forms will likely coexist in any foreseeable future.
A summary of the paper’s organization is briefly illustrated in Fig. 1. The paper begins by exploring the necessity for different definitions of intelligence, provide some of the expert opinions on the subject and its origins which later is followed by the introduction of the novel notion of artificial human intelligence, arguing in favor of the inevitable confluence of artificial and natural intelligences. This unique interaction, we hypothesize, would be the defacto standard of future human-to-human as well as human-to-machine communications and interfaces. It must be emphasized though that such an interaction is not an alternative to Artificial General Intelligence (AGI) [4, 5, 6] in any sense, but rather a tentative direction achieving more generic and capable machine intelligence with arguably so many human-like features at behavioral, architectural and functional levels [2]. Later, we provide details about what we mean by artificial human intelligence and review few attempts as the state-of-the-art (SoTA) in this field. We further use brain-centric categorization of past efforts and consider the missing elements in brain-inspired information processing. We delve into critical questions that concern where exactly humans will be positioned in the newly formed artificial intelligence ecosystem, addressing few challenges and discuss the final alignment with humans. Finally, we comment/speculate on the effects of AI on the human brain and cognition and the set of required management actions in the long term as part of our discussions.
2 Views on Intelligence
@@
The quest for defining intelligence is a multifaceted challenge, labyrinthine in its nature and diverse in its interpretations. Intelligence includes a spectrum of cognitive abilities, adaptive behaviors, and problem-solving skills exhibited predominantly by living organisms. At its center, intelligence refers to the capacity to acquire knowledge, reason abstractly (including counterfactual thinking), solve problems, learn from exploratory experience (through interactions with the three dimensional world model), and adapt to unseen circumstances through generalization, quickly without necessarily using many examples (one-shot learning [11]). Intelligence might manifest itself in diverse forms across species, each tailored to suit the ecological niche and evolutionary context of the organism. Human intelligence, often pointed at as the cornerstone of cognitive evolution, is characterized by its complexity, versatility, and capacity for symbolic thought, language, and cultural and societal intercourse. However, defining human intelligence as the ultimate form of intelligence risks anthropocentric bias and overlooks the diverse manifestations of intelligence found in other organisms [12] or more recently in self-learning machines [13].
@@
Scholars have proposed various models and frameworks to elucidate the multidimensional nature of intelligence. Gardner’s theory of multiple intelligences [14] posits that the broad notion of intelligence should comprise distinct modalities, including linguistic, logical-mathematical, spatial, musical, bodily-kinesthetic, inter-personal, intra-personal, and naturalistic intelligences. Each intelligence represents a domain of expertise and proficiency, offering insights into the diverse talents and capabilities of individuals. On the other hand, recent developments in deep learning have enabled rather constrained but various forms of intelligences capable of performing complex tasks and sometimes exhibiting human-like cognitive abilities [15]. This pluralistic account of intelligence allows a rethinking of developmental trajectory with necessary cross modality interactions. However, replicating the full breadth and depth of human intelligence remains elusive, as deep learning systems often excel in narrow domains but struggle with generalization, common-sense reasoning, and context-dependent understanding. Furthermore, ethical considerations surrounding the learning systems raise concerns about the societal impact of new forms of intelligences, including issues of autonomy, accountability, and bias. As we venture into the realms of various forms of artificial and/or non-human intelligence, navigating these challenges requires careful consideration of ethical, social, and philosophical implications.
@@
3 Origins and the Path leading to AHI
@@
Neither the origin of intelligence nor a theory for this origin have been adequately considered and addressed by the past literature. In fact, all definitions of intelligence seem to have a strong subjective component. Whether it be Turing or Newell [22], their descriptions of intelligence, involving knowledge-level considerations, beliefs, and intentions to a system, is strictly based on human judgements making both of them subjective. On the contrary, Steel [23] supports an alternative definition which highlights the continuum nature of it which progressively evolves from chemical systems to actual living organisms. Through precise definition of evolving complex adaptive systems, the quest for origins is reduced down to addressing two fundamental questions, namely, (1) how the brain demonstrates its remarkable behavioral plasticity and (2) how a symbolic layer might have emerged early in the development [23, 24, 25].
Based on our past human experience and historical account that intelligence most likely emerged from active perception and tight/attentive mechanisms that entail 3D interactions with the real-world phenomena through multi-sensory input [26, 27]. In other words, the biological intelligence requires to exist in multi-dimensional feedback-enabled open environments to emerge and develop [27]. We can think of different models of natural intelligence [14] [^2] According to Gardner, natural(istic) intelligence refers to the ability to recognize, categorize, and interact with elements of the environment, including objects, animals, and plants. This type of intelligence enables us to distinguish between different species, groups of people, or objects and to comprehend their interrelationships., but one of the observable models that is immediately accessible and intriguing is the human intelligence which has been leveraged by machine intelligence innovators (particularly deep learning community) in the past for so many years to inspire novel architectures, training methodologies, learning techniques, etc. Multi-model sensory input processing is essential to human intelligence as there is considerable evidence to support their joint progression early in childhood [28, 29]. In fact, recent work on the computational modellings of visually guided haptic sensory systems advocate for the value of multimodal fusion to improve the visual experience and performance [30]. As different sensory systems can be linked to distinct intelligent behaviors, growing number of studies seem to suggest that their synchronization and fusion lead to more accuracy in completion of specific tasks [31], which demonstrates the potential for integrating more general and non-biological intelligences. One of the most compelling pieces of evidence supporting this finding is the success of multimodal foundation models, which have demonstrated that a missing modality can be compensated for by the presence of other modalities as part of their emergent behavior [32].
Inspired from the past studies, we characterize AHI to describe trichotomous, consisting of human-assisted, human-inspired or human-independent machine intelligences. Since each one of these trejectories are rooted based on human capabilities and human intelligence, we coined the name AHI to describe all. Just like found in the integration of multimodal sensory inputs, this paradigm’s objective is to improve system performance by leveraging the distinct learning mechanisms of both humans and various AI approaches, which might result in complementarity in many aspects of processing and eventually in the final task performance. Additionally, the involvement of humans or understanding tasks performed by the human brain can help regulate ethical considerations, symbolic computations, and emotional aspects of the underlying functional systems.
@@
Figure 2: In human-inspired intelligence, developments in machine intelligence are primarily driven by insights from biology and neuroscience, leading to overlapping advancements in both fields as each one of them evolves over time. Conversely, in human-assisted intelligence, the overlap tends to diminish over time, resulting in more specialized assistance from either humans or machines, thereby fostering the development of hybrid systems. Finally, in human-independent intelligence, the intrinsic features of machine intelligence evolve freely, naturally causing a divergence in the fundamental principles governing human and machine intelligence as the constraints such as hardware that hold up these intelligences become more diverse.
3.1 The current SoTA and Taxonomy
@@
Current AI systems do not allocate enough emphasis to the definition of the intelligence problem, nor they seem to necessarily align in their objectives with different forms of human and/or non-human intelligences. Most early literature was inclined to adopt the model of human intelligence from a pragmatic perspective: The layered models of mind [33] (see also AlexNet [34] or VGGNet [35] etc.), a representation of layered cortical structures and hierarchical activations of neural populations, inspired inductive biases which eventually got encoded into the model architectures. They typically define the elevation to this intelligence as getting the job done faster, more efficient and accurate than most humans do. While engineering the machines, the evolution of the intelligence (and its definition) will not necessarily be following the same trajectory and hence be compatible with the model of human intelligence and its development. Although many of the most popular architectural models, such as residual networks [36, 37, 38, 39, 40], primarily designed to address a hardware or an algorithm issue, later has been shown to possess biological parallels both physiologically and functionally[15]. There are already many attempts in literature exploring this implicit comparison with respect to human systems [41, 42, 43].
@@
3.2 Human-inspired, Human-assisted, Human-Independent AI
The distinction between “human-inspired” and “human-assisted” AI pertains to the extent of human involvement and influence in the development and operation of artificial intelligence systems. A summary of this taxonomy is illustrated geometrically in Fig. 2, along with the temporal progression (directed arrows) of each paradigm in the evolutionary trajectory of the work that has been done in the last few decades (from Past to Future). More specifically, assuming stable centroids of the studies conducted around human and machine intelligences, a human-inspired trajectory will expand future works around these centroids, thereby naturally leading to more commonalities, i.e., increased biological plausibility due to everlasting incremental research around the same or similar ideas. In the human-assisted approach, the trend is reverse i.e., the future overlap between machine and human intelligences tend to shrink allowing hybrid designs to emerge and have better capacity to effectively address intelligence problem. Human assistance is not used as part of the training processs but rather used as the essential component of the system. On the contrary, human-independent trend moves the centroids of the literature works in different directions while their expansion is likely to continue, potentially creating more overlaps between the two distinct intelligences. Fig. 3 illustrates these options in detail and in the following subsections, we shall closely explore verbally each one of these options.
3.2.1 Human–inspired AI: Neuroscience meets Machines
In Human-inspired AI, machines are designed to mimic or replicate aspects of human intelligence, behavior, or complex cognitive processes. Part of this endevour could be inspired by the brain mechanisms responsible for processing information and navigating attention to complete complex tasks. Human-inspired AI aims at understanding and replicating the neural mechanisms behind human cognitive processes, such as learning, common sense reasoning, perception, and critical decision-making. This approach often involves studying biological systems closely including neuroscience, psychology, cellular chemistry and cognitive sciences to accurately inform and guide the design of AI algorithms and architectures. Perhaps the most critical phase of the formulation of human–like intelligence is the developmental phases of the brain anatomy and its training (learning trajectory) with psychophysical limitations impacting its evolution. Examples include neural networks inspired by the brain’s architecture (structural and functional connectivity) and specialization of the specific human brain regions, biologically plausible learning algorithms [44] or algorithms (recursive, top-down and bottom-up) that accurately model human decision-making processes. Human-inspired AI can be designed through methodological changes in model architecture or training. Architectural changes could be implemented through structural or functional changes involving addition or removal of nuronal units in the process of training/learning processes.

### v2

@@ Abstract
Human intelligence, the most evident and accessible form of source of reasoning, hosted by biological hardware, has evolved and been refined over thousands of years, positioning itself today to create new artificial forms and preparing to self–design their evolutionary path forward. Beginning with the advent of foundation models, the rate at which human and artificial intelligence interact with each other has exceeded any anticipated quantitative figures. The close engagement led both bits of intelligence to be impacted in various ways, which naturally resulted in complex confluences that warrant close scrutiny. In the sequel, using a novel taxonomy, we shall explore the interplay between human and machine intelligence, focusing on the crucial role humans play in developing ethical, responsible, and robust intelligent systems. We briefly delve into various aspects of implementation inspired by the mechanisms underlying neuroscience and human cognition. In addition, we propose future perspectives, capitalizing on the advantages of symbiotic designs to suggest a human-centered direction for next-generation developments, focusing on the augmentation role of AI. We finalize this evolving document with some thoughts and open questions yet to be addressed by the broader community.
Index Terms:
Human-level AI, neuroscience, vision, cognition, neural networks, biological plausibility, CNN.
I Introduction
@@
The complexity of the intelligence problem, to date still resisting a universally accepted definition, often motivates adaptation of the human model in identifying intelligent behavior. Artificial intelligence, in most forms conceivable, traces its origin with either mere mimicking of human functionality or a direct product of human intellect, inspired from various sources such as outside physical world. This dependence inherently biases its design. The taxonomy briefly illustrated in Fig. 1 reveals various sources of inspiration for human-level AI development, ranging from bioinspired and brain-inspired to purely symbolic algorithmic constraints. The same illustration clearly marks the definition of the novel concept of Artificial Human Intelligence (AHI) that spans all layers of inspiration and impacts the design evolution of next-generation AI around human-centric skills. Human-inspired intelligence, as a subclass of AHI, acts as a critical category that will actually interlink human cognition with a wide variety of machine capabilities. Fostering collaboration among these domains could enable humankind to harness revolutionary cognitive technologies and inventions that push the boundaries of human thought. Moving through multiple layers of inspiration, it forms a synergistic feedback loop that will yield high human-level AI, representing the emerging relationship between digital systems and biological computation models.
In this brief exposition, our objective is to investigate the strategic placement of human–centered inquiries in the advancement of contemporary AI systems despite the recent remarkable (not necessarily human–centric) developments in machine learning and computer vision fields [2]. Considering the strategies of machine and human intelligences in completing various tasks and their complementarity in behavior, a potential fusion of both emerges as a plausible research direction that might soon become an applicable reality. Due to the involvement of humans as biological organisms, more specifically brain-centric investigations, to be implemented on pure algorithm-driven computational platforms, we have decided to label this specific fusion of intelligences and all the interactions within as AHI and explicate various intriguing topics of discourse for the mutually beneficial association of the two (seemingly) mechanistically distinct notions [3].
Figure 1: The kinds of inspiration sources in search of modeling intelligence with their corresponding levels of complexity and explainability within the context of Artificial Human Intelligence, a concept that bridges (through inspiration) between brain-inspired and human-independent intelligences.
Naturally, the exploration of both forms of intelligence necessitates addressing numerous intriguing questions. These include early human brain developments and their connections to neuroscience, as well as biological hardware dependencies, efficiency of implementations and the potential long-term effects on both human and machine intelligences as both intelligent forms will likely coexist in any foreseeable future.
A summary of the paper’s organization is briefly illustrated in Fig. 2. Main objective of the paper, role of Humans in the development of next gen-AI, is introduced in Section I. By exploring the necessity for different definitions of intelligence, the following section (Section II) provide some of the expert opinions on the subject. In Section III, we discuss the origins of biological intelligence later followed by the introduction of the novel notion of artificial human intelligence in Section III-B, arguing in favor of the inevitable confluence of artificial and natural intelligences (see Fig. 1). This unique interaction, we hypothesize, would be the de facto standard of future human-to-human as well as human-to-machine communications and interfaces. It must be emphasized though that such an interaction is not an alternative to Artificial General Intelligence (AGI) [4, 5, 6] in any sense, but rather a tentative direction achieving more generic and capable machine intelligence with arguably so many human-like features at behavioral, architectural and functional levels [7]. In Section IV, we delve into human-level AI, some of the questions that concern where exactly humans will be positioned in the newly formed AI ecosystem, addressing few challenges and discuss the final alignment with humans. Finally in Section V, we comment/speculate on the effects of AI on the human brain and cognition and the set of required management actions in the long term as part of our discussions, before we conclude the paper in Section VI.
II Views on Intelligence
@@
The quest for defining intelligence is a multifaceted challenge, labyrinthine in its nature and diverse in its interpretations. Intelligence, broadly conceived, includes a spectrum of cognitive abilities, adaptive behaviors, and problem-solving skills exhibited predominantly by living organisms. At its core, intelligence involves the capacity to acquire knowledge, reason abstractly (including counterfactual thinking), solve problems, learn from exploratory experience (through interactions with the three dimensional world model), and adapt to unseen circumstances (generalization) quickly without necessarily using many examples (one-shot learning [12]). Intelligence might manifest itself in diverse forms across species, each tailored to suit the ecological niche and evolutionary context of the organism. Human intelligence, often hailed as the pinnacle of cognitive evolution, is characterized by its complexity, versatility, and capacity for symbolic thought, language, and cultural and societal intercourse. However, defining human intelligence as the ultimate form of intelligence risks anthropocentric bias and overlooks the diverse manifestations of intelligence found in other organisms [13] or more recently in self-learning machines [14].
@@
Scholars have proposed various models and frameworks to elucidate the multidimensional nature of intelligence. Gardner’s theory of multiple intelligences [15] posits that the broad notion of intelligence should comprise distinct modalities, including linguistic, logical-mathematical, spatial, musical, bodily-kinesthetic, inter-personal, intra-personal, and naturalistic intelligences. Each intelligence represents a domain of expertise and proficiency, offering insights into the diverse talents and capabilities of individuals. On the other hand, recent developments in deep learning have enabled rather constrained but various forms of intelligences capable of performing complex tasks and sometimes exhibiting human-like cognitive abilities [16]. This pluralistic account of intelligence allows a rethinking of developmental trajectory with necessary cross modality interactions. However, replicating the full breadth and depth of human intelligence remains elusive, as deep learning systems often excel in narrow domains but struggle with generalization, common-sense reasoning, and context-dependent understanding. Furthermore, ethical considerations surrounding the learning systems raise concerns about the societal impact of new forms of intelligences, including issues of autonomy, accountability, and bias. As we venture into the realms of various forms of machine intelligence, navigating these challenges requires careful consideration of ethical, social, and philosophical implications.
@@
III Path leading to AHI
III-A Early Origins
@@
Neither the origin of biological intelligence nor a theory for this origin have been adequately considered and addressed by the past literature. In fact, we primarily notice that all definitions of intelligence leading to origins seem to have a strong subjective component. Whether it be Turing or Newell [23], their descriptions of intelligence, involving knowledge-level considerations, beliefs, and intentions to a system, is strictly based on human judgements making both of them subjective. In addition, these approaches adapts discrete steps of emergence towards complex intelligent behavior, rather than a gradual time course. On the other hand, Steel [24] supports an alternative approach which highlights the continuum nature of it: A construct progressively evolving from chemical systems to actual living organisms. Through close inspection of evolving complex adaptive systems, the quest for biological origins can be reduced down to addressing two fundamental questions, namely, (1) how the brain demonstrates its remarkable behavioral plasticity and (2) how a symbolic layer might have emerged early in the development [24, 25].
@@
Figure 3: In HIns Intelligence, developments in machine intelligence are primarily driven by insights from biology and neuroscience, leading to overlapping advancements in both fields as each one of them evolves over time. Conversely, in HAss intelligence, the overlap tends to diminish over time, resulting in more specialized assistance from either humans or machines, thereby fostering the development of hybrid systems. Finally, in HInd intelligence, the intrinsic features of machine intelligence evolve freely, naturally causing a divergence in the fundamental principles governing human and machine intelligence as the constraints such as hardware that hold up these intelligences become more diverse.
Based on our past human experience and the related historical account, biological intelligence most likely emerged from active perception and tight/attentive mechanisms that entail 3D interactions with the real-world phenomena through multi-sensory input [26, 27]. In other words, the biological intelligence requires to exist in multi-dimensional feedback-enabled open environments to emerge and develop [27], which eventually manifests agency-like behaviour. The design of intelligent machines can potentially provide clues about origins as they might likely simulate the emergence. While developing such models, we can think of different models (levels) of natural intelligence [^2] According to Gardner, natural(istic) intelligence refers to the ability to recognize, categorize, and interact with elements of the environment, including objects, animals, and plants. This type of intelligence enables us to distinguish between different species, groups of people, or objects and to comprehend their interrelationships. [15], but one of the observable models immediately accessible and intriguing is the human intelligence which has been heavily leveraged by machine intelligence innovators (particularly deep learning community) in the past for so many years to inspire novel architectures, training methodologies, learning techniques, etc. Multi-model sensory input processing is essential to human intelligence as there is considerable evidence to support their joint progression early in childhood [28, 29]. In fact, recent work on the computational modelings of visually guided haptic sensory systems advocate for the value of multimodal fusion to improve the visual experience and performance [30]. As different sensory systems can be linked to distinct intelligent behaviors, growing number of studies seem to suggest that their synchronization and fusion lead to improved accuracy in completion of specific tasks [31], which demonstrates the potential for integrating more general and non-biological intelligences. One of the most compelling pieces of evidence supporting this finding is the success of multimodal foundation models, which have demonstrated that a missing modality can be compensated for by the presence of other modalities [32].
Inspired from the past studies, we characterize AHI to describe trichotomous, consisting of the following three kinds of machine intelligences,
  • Human-Inspired (HIns): AI systems that draw inspiration from human cognitive processes, such as learning, reasoning, or problem-solving, to mimic human-like decision-making and understanding.
  • Human-Assisted (HAss): AI systems that rely on human input, collaboration, or feedback for operation, training, or improvement, often combining human expertise with machine efficiency.
  • Human-Independent (HInd): Autonomous self-improving AI systems capable of operating, learning, and making decisions without direct human involvement, designed to perform tasks independently.
Just like found in the integration of multimodal sensory inputs, this paradigm’s objective is to improve system performance by leveraging the distinct learning mechanisms of both humans and various AI approaches, which might result in complementarity in many aspects of processing and eventually in the final task performance. Additionally, the involvement of humans or understanding tasks performed by the human brain can help regulate ethical considerations, symbolic computations, and emotional aspects of the underlying functional systems.
III-B The current SoTA and Taxonomy
@@
Current AI systems do not allocate enough emphasis to the definition/origins of the intelligence problem, nor they seem to necessarily align in their objectives with different forms of human or machine intelligences. Most early literature was inclined to adopt the model of human intelligence from a pragmatic perspective: The layered models of mind [33] (see also AlexNet [34] or VGGNet [35] etc.), a representation of layered cortical structures and hierarchical activations of neural populations, inspired inductive biases which eventually got encoded into the model architectures. They typically define the elevation to this intelligence as getting the job done faster, more efficient and accurate than most humans do. While engineering the machines, the evolution of the intelligence (and its definition) will not necessarily be following the same trajectory and hence be compatible with the model of human intelligence and its development. Although many of the most popular architectural models, such as residual networks [36, 37, 38, 39, 40], primarily designed to address a hardware or an algorithm issue, later has been shown to possess biological parallels both physiologically and functionally[16]. There are already many attempts in literature exploring this implicit comparison with respect to human systems [41, 42, 43].
@@
III-C Comparisons between HIns, HAss and HInd Intelligences
The distinction between “human-inspired” and “human-assisted” AI pertains to the extent of human involvement and influence in the development and operation of artificial intelligence systems. A summary of this taxonomy is illustrated geometrically in Fig. 3, along with the temporal progression (directed arrows) of each paradigm in the evolutionary trajectory of the work that has been done in the last few decades (from Past to Future). More specifically, assuming stable centroids of the studies conducted around human and machine intelligences, a human-inspired (HIns) trajectory will expand future works around these centroids, thereby naturally leading to more commonalities, i.e., increased biological plausibility due to everlasting incremental research around the same or similar ideas. In the human-assisted (HIns) approach, the trend is reverse i.e., the future overlap between machine and human intelligences tend to shrink allowing hybrid designs to emerge and have better capacity to effectively address intelligence problem. Human assistance is not used as part of the training process but rather used as the essential component of the system. On the contrary, human-independent (HInd) trend moves the centroids of the literature works in different directions while their expansion is likely to continue, potentially creating more overlaps between the two distinct intelligences. Fig. 4 illustrates these options in detail and in the following subsections, we shall closely explore verbally each one of these options along with the associated subcategories.